# Agent-as-Judge for Factual Summarization of Long Narratives

### v2

@@ 5.4 Latency
Table 7: Average latency (in minutes) and KENDALL’s tau correlation for evaluating factuality across different metrics on the FABLES Kim et al. (2024) dataset.
@@ 5.5 Sensitivity of Metrics on Factual Perturbation
We evaluated metric sensitivity to factual perturbations using GPT-4o with the prompt shown in Figure 10 on the MENSA. Specifically, we perturbed the reference summaries from MENSA by introducing factual inaccuracies in each sentence. Table 8 shows that ROUGE-L and BERTScoref1 decreased minimally despite factual perturbations, while NarrativeFactScore significantly dropped. These results demonstrate that our metric is highly sensitive to factual discrepancies, making it a suitable metric for assessing factuality.
@@ 6 Conclusion
This work shows how the agent-as-judge is deployed to overcome the limitations of existing evaluation metrics, such as overreliance on lexical similarity or factual inconsistencies. Specifically, we propose consistent CKG extraction, and new factual evaluation metric based on CKG, and an agent that evaluates and guides the summary and refinement. Through our implementation, we demonstrated both the process and superior performance over state-of-the-art methods on real-life industry datasets and scenarios.
@@ 7 Limitation
Appendix A Qualitative Example
@@
Appendix B System Deployment
@@
This section describes our system deployment, which is necessary for the media industry [^10] Company name anonymized for blind reviewing. The media industry broadly refers to the sector that creates, distributes, and analyzes various forms of narratives, such as movies, television shows, books, video games, and other media that tell stories. This includes businesses involved in producing, editing, and consuming these forms of content, focusing on storytelling in both traditional and digital media. where companies make investment decisions on narratives (> 100K tokens) such as dramas or movies. Since reading all narratives is challenging, media companies utilize summaries of each narrative to determine its production feasibility. However, summaries generated by humans or LLMs frequently contain factual inconsistencies, which hinder accurate investment decisions. Therefore, our proposed system is deployed to evaluate the factuality of summary for long narratives and improve its factuality.
Figure 4 shows screenshots of the system111111bit.ly/4iz8pXc, aligned with the three phases of our framework in Figure 2. Using the example Black Panther, users can view the original narrative after selecting a dataset, data type, and name. Clicking “Generate Knowledge Graph” generates and visualizes the CKG (Section 3.1). The “Generate Initial Summary” and “Calculate Factuality Score” buttons create an initial summary and evaluate its factuality using the CKG (Section 3.2). Finally, “Refine Summary” improves the summary based on feedback, enhancing factuality (Section 3.3).
Appendix C Usage of AI Assistants
We utilized ChatGPT to improve the clarity and grammatical accuracy of my writing. It provided suggestions for rephrasing sentences and correcting grammatical errors to make the text flow more naturally.
Appendix D Prompts
@@
D.1 Knowledge Extraction Prompt for LLM
@@
D.2 Narrative Summarization Prompt for LLM
@@
D.3 Atomic Fact Decomposition Prompt for LLM
@@
D.4 Fact-Checking Prompt for NarrativeFactScore
@@
D.5 Agent-based Refinement Prompt for LLM
@@
D.6 Factual Perturbation Prompt

### v3

@@ 5.3 Latency
Table 7: Average latency (in minutes) per a summary and Kendall’s tau correlation for evaluating factuality across different metrics on the FABLES Kim et al. (2024) dataset. Reported total times are the sum of CKG extraction and factuality calculation.
@@
To further analyze efficiency, we also evaluated a one-shot variant of NarrativeFactScore (NFS w/o consistency). As reported in Table 7, this variant requires less time for CKG extraction but shows weaker correlation with human judgments, highlighting a trade-off between speed and accuracy.
@@ 5.4 Sensitivity of Metrics on Factual Perturbation
We evaluated metric sensitivity to factual perturbations using GPT-4o with the prompt shown in Figure 10 on the MENSA. Specifically, we perturbed the reference summaries from MENSA by introducing factual inaccuracies in each sentence. Table 8 shows that ROUGE-L and BERTScoref1 decreased minimally despite factual perturbations, while NarrativeFactScore significantly dropped. In our setup, perturbations were created by minimally replacing specific words in the reference summaries rather than generating entirely new content, since the latter would unfairly penalize ROUGE and BERTScore. These results demonstrate that our metric is highly sensitive to factual discrepancies, making it a suitable metric for assessing factuality.
@@ 6 Conclusion
This work shows how the agent-as-judge is deployed for narrative summarization to overcome the limitations of existing evaluation metrics, such as overreliance on lexical similarity or factual inconsistencies. Specifically, we propose consistent CKG extraction, and a new factual evaluation metric based on CKG, and an agent that evaluates and guides the summary and refinement. Through our implementation, we demonstrated both the process and superior performance over state-of-the-art methods on real-life industry datasets and scenarios.
@@ 7 Limitation
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00414981), the MSIT (Ministry of Science and ICT), Korea, under the ITRC (Information Technology Research Center) support program (IITP-2025-2020-0-01789) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation), and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)].
@@ Acknowledgements
Appendix A Further Analysis
A.1 Analysis for Baseline Metrics
Strength Spearman ($\rho$) KENDALL ($\tau$)
Very weak correlation 0.00~0.15 0.00~0.10
Weak correlation 0.15~0.30 0.10~0.20
Moderate correlation 0.30~0.43 0.20~0.30
Strong correlation 0.43~1.00 0.30~1.00
Table 9: Correlation strength based on coefficients.
[…]
A.2 Effectiveness with Open-Source LLM
R-1 R-2 R-L BSp BSr BSf1 NFS
without merging
     TextRank 34.37 4.60 12.84 46.86 49.43 48.10 59.72
     LED 17.46 1.59 10.03 42.90 42.74 42.58 56.48
     LongT5 20.77 2.26 10.03 45.05 45.06 45.01 73.76
hierarchically merging
     Qwen2.5-14B-Instruct 30.45 7.03 11.99 58.20 59.61 58.89 67.29
     Ours: 1st iteration 31.26 7.06 12.05 58.31 59.68 58.95 70.71
     Ours: 2nd iteration 31.57 7.08 12.09 58.35 59.69 58.97 71.89
Table 10: Evaluation results on the MENSA Saxena and Keller (2024b) dataset. Hierarchically merging results are based on Qwen2.5-14B-Instruct.
Our framework is also effective with open-source model. As shown in Table 10, NarrativeFactScore implemented using Qwen2.5-14B-Instruct Qwen Team (2024) consistently improves summarization performance on MENSA across two iterations. Compared to the initial summarization, our refinement shows steady gains, achieving NFS of 70.71 and 71.89 after the first and second iterations, respectively. These results confirm that the benefits of our method are not limited to proprietary LLMs but also extend to accessible open-source alternatives.
A.3 Baseline Evaluation with Hierarchical Merging
MENSA MovieSum
R-1 R-2 R-L BSp BSr BSf1 NFS R-1 R-2 R-L BSp BSr BSf1 NFS
hierarchically merging
     LED 22.02 3.30 9.00 43.10 43.62 43.19 59.71 21.45 2.55 14.37 44.86 44.32 44.49 61.90
     LongT5 30.88 4.49 11.23 45.85 48.84 47.27 78.13 31.19 4.06 17.62 46.01 48.92 47.39 79.15
Table 11: Evaluation results on MENSA and MovieSum datasets with hierarchical merging.
Our conclusions remain consistent even when baselines are evaluated with hierarchical merging. As shown in Table 11, NarrativeFactScore with refinement continues to outperform LED and LongT5 on both MENSA and MovieSum, confirming the robustness of our improvements regardless of baseline settings.
Appendix B Qualitative Example
@@
Initial
summary
Back at Bag End, Gandalf confronts Frodo about the Ring, revealing its dark history and the resurgence of Sauron, who seeks it to regain power.
Atomic
facts
[1] Gandalf confronts Frodo about the Ring at Bag End.
[2] Gandalf reveals the Ring’s dark history to Frodo.
[3] Sauron seeks the Ring to regain power.
Factuality
calculation
[1] True
[2] True
[3] False, The statement is judged false because the retrieved subgraph includes the triple “Sauron-master of Ring”, implying that Sauron already possesses the Ring and does not need to seek it.
(evidence scene: #29, evidence subgraph: ‘Sauron-master-Ring’, …)
Refined
summary
Back at Bag End, Gandalf confronts Frodo about the Ring, revealing its dark history and the resurgence of Sauron, who already possesses the Ring and does not seek it.
Atomic
facts
[1] Gandalf confronts Frodo about the Ring at Bag End.
[2] Gandalf reveals the Ring’s dark history to Frodo.
[3] Sauron already possesses the Ring and does not seek it.
Factuality
calculation
[1] True
[2] True
[3] True, Correction: The retrieved subgraph was irrelevant. In fact, Frodo is the Ring-bearer, and Sauron is seeking the Ring to regain power. Thus, the correct statement is that Sauron seeks the Ring.
(evidence scene: #25, evidence subgraph: ‘Frodo-own-Ring’, ‘Sauron-desire-Ring’, …)
Refined
summary
Back at Bag End, Gandalf confronts Frodo about the Ring, revealing its dark history and the resurgence of Sauron, who seeks it to regain power.
Table 13: Qualitative example illustrating NarrativeFactScore when irrelevant subgraph is retrieved. Blue text in the initial summary and atomic facts indicates the correct fact, but due to irrelevant retrieval, it is first misjudged as false (red). Through subsequent refinement, the error is corrected.
As shown in Table 13, this example shows how irrelevant subgraph retrieval can cause errors. The atomic fact [3] “Sauron seeks the Ring to regain power” is correct. However, the retrieved triple “Sauron-master of Ring” gives misleading feedback that Sauron already owns the Ring. This results in an incorrect refinement, where the fact is wrongly judged as false. Through iterative refinement, the system retrieves more relevant subgraphs (e.g., Frodo owns the Ring and Sauron seeks it) or uses scene evidence. As a result, the correct factual judgment is recovered in later iterations.
Appendix C System Deployment
@@
This section describes our system deployment, which is necessary for the media industry [^10] The media industry broadly refers to the sector that creates, distributes, and analyzes various forms of narratives, such as movies, television shows, books, video games, and other media that tell stories. This includes businesses involved in producing, editing, and consuming these forms of content, focusing on storytelling in both traditional and digital media. where companies make investment decisions on narratives (> 100K tokens) such as dramas or movies. Since reading all narratives is challenging, media companies utilize summaries of each narrative to determine its production feasibility. However, summaries generated by humans or LLMs frequently contain factual inconsistencies, which hinder accurate investment decisions. Therefore, our proposed system is deployed to evaluate the factuality of summary for long narratives and improve its factuality.
Figure 4 shows screenshots of the system111111huggingface.co/spaces/yeonseokjeong/NarrativeFactScore, aligned with the three phases of our framework in Figure 2. Using the example Black Panther, users can view the original narrative after selecting a dataset, data type, and name. Clicking “Generate Knowledge Graph” generates and visualizes the CKG (Section 3.1). The “Generate Initial Summary” and “Calculate Factuality Score” buttons create an initial summary and evaluate its factuality using the CKG (Section 3.2). Finally, “Refine Summary” improves the summary based on feedback, enhancing factuality (Section 3.3).
Appendix D Usage of AI Assistants
We utilized ChatGPT to improve the clarity and grammatical accuracy of the writing. It provided suggestions for rephrasing sentences and correcting grammatical errors to make the text flow more naturally.
Appendix E Prompts
@@
E.1 Knowledge Extraction Prompt for LLM
@@
E.2 Narrative Summarization Prompt for LLM
@@
E.3 Atomic Fact Decomposition Prompt for LLM
@@
E.4 Fact-Checking Prompt for NarrativeFactScore
@@
E.5 Agent-based Refinement Prompt for LLM
@@
E.6 Factual Perturbation Prompt